OpenStudy (anonymous):

complete the square y^2+36=13y

OpenStudy (anonymous):

y^2+36=13y y^2- 13y + 36=0 y^2- 13y = - 36 -13/2 y^2 -13y + (-13/2)^2= -36 +(-13/2)^2 can you finish it

OpenStudy (anonymous):

y^2+(-13/2)^2=49+(-13/2)^2 is tht right so far?

OpenStudy (anonymous):

13y can NOT be added or subtracted from 36

OpenStudy (anonymous):

ohh haha i totally didn't think about tht :D

OpenStudy (anonymous):

i will do it using letters y^2+b=ay y^2-ay=-b take 1/2 *a then square it (a/2)^2 lets make c= (a/2)^2 y^2 - ay + c =-b + c (y-a)^2 = -b + c try it

OpenStudy (anonymous):

huh what happened there

OpenStudy (anonymous):

sorry i got the answer :)